# A Unified Approach for Multi-granularity Search over Spatial Datasets

### v2

@@ III-B2 Point-wise Similarity Measure
III-C Types of Dataset Search
@@
(Top-$k$ Exemplar Dataset Search (ExempS)). Given a query dataset $Q$ and a spatial data repository $\mathcal{D}$, a top-$k$ exemplar search returns $k$ datasets from $\mathcal{D}$ that are most similar to $Q$, i.e., it retrieves $\mathcal{D}_{q}\subseteq\mathcal{D}$ with $k$ datasets such that: $\forall D\in\mathcal{D}_{q}$ and $\forall D^{\prime}\in\mathcal{D}\backslash\mathcal{D}_{q}$, we have $dist(Q,D)\leq dist(Q,D^{\prime})$, where $dist$ is any of the distance measures defined in Section III-B.
III-D Types of Data Point Search
@@ III-E Problem Formulation
RoadMap.  The rest of this paper is organized as follows. We first introduce the overall system architecture in Section IV. Then, we describe the unified index and how it is constructed in Section V. Subsequently, we introduce a series of effective acceleration and pruning mechanisms in Section VI. Finally, experimental results are provided in Section VII.
IV Spatial Data Search System Framework
Given the problem described in Section III, we propose a multi-granularity spatial data search system called Spadas. The architecture of Spadas and its components are shown in Fig 3. It consists of two layers that we describe below.
Figure 3: Architecture of Spadas.
IV-A Index Layer
Existing indexes mainly adopt a single-level index structure that only focuses on dataset-level or data point-level searches. To effectively support a rich set of queries presented in Section III, we design a novel two-level unified index structure in the index layer, including the upper-level index and bottom-level index, effectively organizing datasets and data points.
@@ IV-A Index Layer
Upper-level Index.  The upper-level index is designed to partition and organize datasets at a higher level. Specifically, to avoid scanning all datasets and achieve a fast dataset search for a given query, we pre-organize the indexes of all datasets using a fast space-splitting indexing mechanism, such as the ball-tree [48]. This allows us to construct an upper-level index, referred to as the data repository index, which bounds all the dataset nodes. The index construction details are elaborated in Section V.
IV-B Search Layer
@@
Pruning Mechanisms.  Based on building a two-level index, we design a set of unified pruning mechanisms to accelerate multiple types of search operations without constructing additional indexes. Specifically, when the query region $R$ or the query dataset $Q$ arrives, we traverse the index from the upper-level index to the bottom-level index in a depth-first manner. Then, we take the branch-and-bound strategy to batch prune out those nodes that cannot satisfy the query request, significantly speeding up the dataset search. The detailed pruning mechanisms are elaborated in Section VI.
V Optimization of Index Layer
Figure 4: An overview of our unified index.
This section provides a detailed introduction to the index layer. Specifically, we design a unified index based on the whole spatial data repository (see Def. 3), which includes a bottom-level index (see Section V-A) and an upper-level index (see Section V-B). Figure 4 presents an overview of our indexing framework. The index is built over a tree structure with nodes bounded by balls and boxes, and we store such bounding information to support multiple types of queries on spatial data. Algorithm 1 shows the details about the implementation of index construction, which will be further explained in the following sections. In the subsequent subsections, we will introduce how the bottom-level (Section V-A) and upper-level index (Section V-B) are constructed in detail.
@@ 59            RefineBottomUp($N_{l}$, $r{{}^{\prime}}$);
V-A Bottom Level: Index of Each Single Dataset
@@
Step2: Parameter-free Outlier Removal.  Most existing methods utilize the dataset index to remove outliers that have a limited number of neighbors within a threshold distance, and reconstruct the index [32, 61, 60, 49, 15]. Their common drawback is the high complexity and the requirement of specifying parameters such as the number of neighbors and threshold distance. Thus, we design a parameter-free mechanism to efficiently remove outliers in spatial datasets, utilizing the distribution of leaf node sizes in the dataset index (see Function OutlierRemoval of Algorithm 1).
We observe in Figure 5 that there is only a small number of leaf nodes with a large radius in the index built for the two data repositories considered (see Section VII for more details). We infer that the leaf nodes with large radius may contain outliers because outliers are normally far from other points in the datasets. Hence, a key question here is how to determine an appropriate radius to distinguish between nodes with or without outliers.
The main idea of our method is illustrated in Figure 6. Firstly, it computes the radii of all leaf nodes and sorts them in an array $\phi$. It then extends the Kneedle algorithm [56, 6] to automatically select an appropriate threshold $r^{\prime}$ based on the sorted array $\phi$. Specifically, it computes a gap ($g_{i}$) for a radius $\phi[i]$ as (see Lines 1 to 1):
$\small g_{i}=\phi[0]-i\times\frac{\phi[0]-\phi[|\phi|-1]}{|\phi|}-\phi[i],i\in\{1,|\phi-1|\}.$ (3)
@@
Figure 5: The distributions of radii of leaf nodes before and after outlier removal on two data repositories.
Figure 6: An illustration of the proposed parameter-free outlier removal method based on the Kneedle algorithm.
@@
V-B Upper Level: Data Repository Index
@@
Time Complexity.  Additionally, we provide a theoretical analysis of the time complexity for constructing the unified index, which is $O(mn\log\left(n\right))$, where $m$ and $n$ represent the number of data points in the dataset and the number of datasets in the data repository, respectively. The detailed analysis is provided in Appendix IX-B.
@@
VI Optimization of Search Layer
Based on the unified index constructed in the previous section, we propose effective pruning mechanisms to solve the data search problem defined in Section III-E.
VI-A Accelerating Dataset Search
In this subsection, we describe the dataset search algorithm applied to range-based dataset search and top-$k$ exemplar dataset search defined in Section III-C. Algorithm 2 shows the details of the implementation of the algorithm.
VI-A1 Accelerating Range-based Dataset Search
@@ 21                   $dis\leftarrow$ compute $dis$ between $Q$ and $N_{c}$;
VI-A2 Accelerating Top-$k$ Exemplar Dataset Search
@@
Figure 7: The distance bounds between two nodes using bounding boxes (left) and bounding balls (right).
Fast Bound Estimation.  We propose an efficient algorithm based on bounding balls of nodes to estimate the lower and upper bounds of the Hausdorff distance between nodes, which is not affected by the dimensions and only needs a single distance computation between center points. In contrast to the R-tree based incremental approach [46], which requires enumerating all pairs of corners of nodes’ bounding boxes (as illustrated by the four black dotted lines in Figure 7(a)), our method avoids quartic Euclidean distance computations [46] and only requires a single computation.
@@
$\small\left[\max(||o_{1},o_{2}||-r_{2},0),\sqrt{||o_{1},o_{2}||^{2}+r_{2}^{2}}+r_{1}\right].$ (4)
@@ 34            $minUB\leftarrow\min(minUB,\textsf{UpperBound}(N_{Q},N_{C}))$;
Exact Hausdorff.  The above bounds can also be directly used in the fast pairwise Hausdorff distance computation. The main idea is to rank candidates based on two priority queues. Algorithm 3 shows how to prune distance computations based on bounds and achieve early termination without computing the full distance between all pairs of points from $Q$ and $D$. Similar to the incremental traversal (Please refer to Algorithm 4 in [46]) [^2] Readers can refer to Algorithm 4 of [46] for more details. Here, we will skip them as the estimation of the bounds is our main focus., since the goal of Hausdorff is to calculate the maximum nearest neighbor distance, we first create a priority queue in descending order, ranking the distance from the point to their nearest neighbors. Then to compute the minimum distances to the nearest neighbors for points in $Q$, we further create an ascending priority queue to traverse the dataset from close to far. Subsequently, we traverse the index of $Q$ and $D$ from the root nodes, estimate the bounds between the polled pair of nodes, and conduct pruning. The whole algorithm terminates when the first pair of points are polled from the main priority queue. For convenience, we refer to this method as “ExactAppro” in Section VII.
Approximate Hausdorff.  Computing the exact Hausdorff distance can be time-consuming by going to the lowest point level of the dataset index. Instead, we can calculate an error-bounded Hausdorff distance in a more efficient way based on the unified index. In the exact distance computation, we terminate until a pair of points show up from the queue. To avoid such intensive distance computations, we set a rule that once a pair of “small” nodes have a radius smaller than a given threshold $\epsilon$ (see Line 3), we compute their Euclidean distance between their center points. Then we return the distance $||o_{1},o_{2}||$ as the approximate distance, and do not enqueue their child nodes or points anymore. For convenience, we refer to this method as “ApproHaus” in Section VII. Next, we prove that the ApproHaus method in Lemma 1 achieves a guaranteed precision of $2\epsilon$, where $\epsilon$ is the error threshold defined by the user.
@@
$\epsilon=\frac{b^{\uparrow}[0]-b^{\downarrow}[0]}{2^{\theta}}.$ (8)
VI-B Accelerating Data Point Search
VI-B1 Range-based Point Search
@@
VI-B2 Nearest Neighbor Point Search
@@
Novelty of Acceleration Strategies.  According to the links between these technological developments, we can see that the proposed acceleration strategies based on our two-level unified index can solve multiple types of search problems at the same time compared to previous methods, which can only solve a single search problem [40, 35, 14]. In addition, we theoretically derive Hausdorff’s upper and lower bounds using the triangle inequality, and we explore the approximate Hausdorff value in depth with a guaranteed theoretical error.
VII Experiments
We conduct experiments on six real-world dataset repositories to demonstrate the effectiveness and efficiency of Spadas. In Section VII-A, we introduce the experimental setups. Then, we present the performance of our proposed unified index in Section VII-B. Next, we present the experimental results that vary with several key parameters to demonstrate the effectiveness and efficiency of the dataset search and data point search in Sections VII-C and VII-D, followed by our case study in Section VII-E.
TABLE I: Summary of six spatial data repositories.
@@ VII Experiments
TABLE II: Parameter settings.
@@
VII-A Settings
Datasets.  We conduct experiments on six representative spatial dataset repositories, each of which is downloaded from open-source spatial data portals. Table I presents the scale and dimension of each data repository. Figure 8 shows the distributions of the number of points in each data repository.
  • MultiOpen dataset, a real-world spatial data repository comprising datasets collected from three different open-source platforms: Points of Interests (POIs) of Baidu Maps Open Platform [7], BTAA Geoportal [2] and NYU Spatial Data Repository [4].
  • T-drive dataset [76] contains a collection of real-world trajectory datasets generated by over 33,000 taxis in Beijing;
  • Argoverse self-driving datasets [21] which record five-second driving scenarios with thousands of points;
  • ShapeNet [20], large-scale datasets of shapes represented by 3D computer-aided design (CAD) models of objects;
  • Chicago taxi trips [3] dataset, which has 194,992,061 trips in total. As all the trips are quantized every 15 minutes, we split the whole dataset every 15 minutes to create a set of sub-datasets, and there are 283,375 datasets in total;
  • Porto dataset [5] transformed from 400+ taxis’ trajectories in the city of Porto for one year.
Figure 8: Dataset scale distributions in six data repositories.
Implementation.  We implement Spadas using Java 1.8 and run all our experiments on a 2x Intel Xeon Platinum 8268 24C 205W 2.9GHz Processor with 32G memory. Our code and guidelines for reproducibility can be accessed on the GitHub [10]. We also have an online demonstration system [11] on the Argoverse data repository, which supports the dataset and data point search proposed in this paper, and our proposed top-$k$ EMD search [70] is also available.
@@ VII-A Settings
Parameter Settings.  The settings of parameters are summarized in Table II where the default settings are underlined. Below is the description of each parameter:
@@
  • $R$: It determines the range of query rectangular, where $\epsilon$ is the value obtained when the resolution $\theta$ is the default value $5$ according to the Equation 8 in Section VI-A2.
@@
  • ScanGBO [52]: the baseline comparison method needs to scan the candidate datasets to compute the grid-based overlap for each query dataset (see Algorithm 2 in [52]).
  • ScanHaus [46]: It designs the lower and upper bounds of Hausdorff by using two point sets’ MBR, and combines branch-and-bound to accelerate filter out dissimilar datasets.
  • IncHaus [46]: It speeds up the Hausdorff distance computation by incrementally exploring the index of two point sets.
  • INNE [13, 80]: the isolation-based anomaly detection using nearest-neighbor ensembles for outlier removal.
@@
  • kNN [59]: a nearest neighbor search method using the early break strategy.
Figure 9: Efficiency in all stages on various data repositories.
VII-B An Overview of spadas
We first show the running time of seven main steps in Spadas system, including the unified index construction, dataset search (RangeS, IA, GBO, ExactHaus), and data point search (RangeP, NNP). As shown in Figure 9, we can observe that unified index construction is the most time-consuming step in most cases. However, the time taken to construct the index is still less than one minute. Moreover, the top-$k$ Hausdorff search is the slowest in dataset search compared with the IA and GBO due to its high complexity, but its running time is still tolerable.
Figure 10: Index’s construction time and occupied space with increasing data repository scale $m=\{10^{-3},10^{-2},10^{-1}$, 1}×$\times|\mathcal{D}|$.
Figure 11: Top-$k$ overlap-based queries as $k$ increases.
Figure 12: Top-$k$ overlap-based queries as $f$ increases.
Figure 13: Grid resolution $\theta$’s effects on top-$k$ GBO query.
Figure 14: Top-$k$ exact Hausdorff search with increasing $k$.
Figure 15: Approximate Hausdorff search with increasing threshold $\epsilon$ (decreasing resolution $\theta$ in x-axis accordingly).
Efficiency of the Unified Index.  We compare our unified index with the existing work that constructs an independent index using R-tree [46] in IncHaus. Figure 10 presents the construction cost and the occupied space as the data repository scale $m$ increases. It is evident that our proposed unified index outperforms the state-of-the-art method in terms of both index construction time and space footprint. Specifically, with regard to index construction time, the unified index can achieve at most 36.8% speedup over the index of IncHaus. Additionally, in terms of index space, the unified index is reduced by up to 90% over the IncHaus.
(a) Accuracy of Top-$k$ Hausdorff Search
(b) Runtime of Top-$k$Hausdorff Search
Figure 16: Accuracy and runtime of three exemplar searches.
(a) Accuracy of Outlier Removal
(b) Runtime of Outlier Removal
Figure 17: Accuracy and runtime of top-$k$ Hausdorff search.
Figure 18: Capacity’s effects on pairwise Hausdorff search.
Figure 19: Index capacity’s effects on top-$k$ Hausdorff search.
(a) Runtime of overlap-based queries
(b) Runtime of point-wise query
Figure 20: Spatial dimension $d$’s effect on the Chicago dataset.
Figure 21: Range-based data point search as $R$ increases.
Figure 22: Nearest neighbor point search as $s$ increases.
Figure 23: Case study on trajectory datasets.
VII-C Efficiency of Dataset Search

### v3

@@ 3.2.2. Point-wise Similarity Measure
3.3. Types of Dataset Search
@@
(Top-$k$ Exemplar Dataset Search (ExempS)). Given a query dataset $Q$ and a spatial data repository $\mathcal{D}$, a top-$k$ exemplar search returns $k$ datasets from $\mathcal{D}$ that are most similar to $Q$, i.e., it retrieves $\mathcal{D}_{q}\subseteq\mathcal{D}$ with $k$ datasets such that: $\forall D\in\mathcal{D}_{q}$ and $\forall D^{\prime}\in\mathcal{D}\backslash\mathcal{D}_{q}$, we have $dist(Q,D)\leq dist(Q,D^{\prime})$, where $dist$ is any of the distance measures defined in Section 3.2.
3.4. Types of Data Point Search
@@ 3.5. Problem Formulation
RoadMap.  The rest of this paper is organized as follows. We first introduce the overall system architecture in Section 4. Then, we describe the unified index and how it is constructed in Section 5. Subsequently, we introduce a series of effective acceleration and pruning mechanisms in Section 6. Finally, experimental results are provided in Section 7.
4. Spatial Data Search System Framework
Given the problem described in Section 3, we propose a multi-granularity spatial data search system called Spadas. The architecture of Spadas and its components are shown in Fig 3. It consists of two layers that we describe below.
Figure 3. Architecture of Spadas.
4.1. Index Layer
Existing indexes mainly adopt a single-level index structure that only focuses on dataset-level or data point-level searches. To effectively support a rich set of queries presented in Section 3, we design a novel two-level unified index structure in the index layer, including the upper-level index and bottom-level index, effectively organizing datasets and data points.
@@ 4.1. Index Layer
Upper-level Index.  The upper-level index is designed to partition and organize datasets at a higher level. Specifically, to avoid scanning all datasets and achieve a fast dataset search for a given query, we pre-organize the indexes of all datasets using a fast space-splitting indexing mechanism, such as the ball-tree (Omohundro1989, ). This allows us to construct an upper-level index, referred to as the data repository index, which bounds all the dataset nodes. The index construction details are elaborated in Section 5.
4.2. Search Layer
@@
Pruning Mechanisms.  Based on building a two-level index, we design a set of unified pruning mechanisms to accelerate multiple types of search operations without constructing additional indexes. Specifically, when the query region $R$ or the query dataset $Q$ arrives, we traverse the index from the upper-level index to the bottom-level index in a depth-first manner. Then, we take the branch-and-bound strategy to batch prune out those nodes that cannot satisfy the query request, significantly speeding up the dataset search. The detailed pruning mechanisms are elaborated in Section 6.
5. Optimization of Index Layer
Figure 4. An overview of our unified index.
This section provides a detailed introduction to the index layer. Specifically, we design a unified index based on the whole spatial data repository (see Def. 3), which includes a bottom-level index (see Section 5.1) and an upper-level index (see Section 5.2). Figure 4 presents an overview of our indexing framework. The index is built over a tree structure with nodes bounded by balls and boxes, and we store such bounding information to support multiple types of queries on spatial data. Algorithm 1 shows the details about the implementation of index construction, which will be further explained in the following sections. In the subsequent subsections, we will introduce how the bottom-level (Section 5.1) and upper-level index (Section 5.2) are constructed in detail.
@@ 59            RefineBottomUp($N_{l}$, $r{{}^{\prime}}$);
5.1. Bottom Level: Index of Each Single Dataset
@@
Step2: Parameter-free Outlier Removal.  Most existing methods utilize the dataset index to remove outliers that have a limited number of neighbors within a threshold distance, and reconstruct the index (Gupta2017, ; Tran2020a, ; Tran2016, ; Orair2010, ; Bhaduri2011, ). Their common drawback is the high complexity and the requirement of specifying parameters such as the number of neighbors and threshold distance. Thus, we design a parameter-free mechanism to efficiently remove outliers in spatial datasets, utilizing the distribution of leaf node sizes in the dataset index (see Function OutlierRemoval of Algorithm 1).
We observe in Figure 5 that there is only a small number of leaf nodes with a large radius in the index built for the two data repositories considered (see Section 7 for more details). We infer that the leaf nodes with large radius may contain outliers because outliers are normally far from other points in the datasets. Hence, a key question here is how to determine an appropriate radius to distinguish between nodes with or without outliers.
The main idea of our method is illustrated in Figure 6. Firstly, it computes the radii of all leaf nodes and sorts them in an array $\phi$. It then extends the Kneedle algorithm (Satopaa2011, ; kneed, ) to automatically select an appropriate threshold $r^{\prime}$ based on the sorted array $\phi$. Specifically, it computes a gap ($g_{i}$) for a radius $\phi[i]$ as (see Lines 1 to 1):
(3) $\small g_{i}=\phi[0]-i\times\frac{\phi[0]-\phi[|\phi|-1]}{|\phi|}-\phi[i],i\in\{1,|\phi-1|\}.$
@@
Figure 5. The distributions of radii of leaf nodes before and after outlier removal on two data repositories.
Figure 6. An illustration of the proposed parameter-free outlier removal method based on the Kneedle algorithm.
@@
5.2. Upper Level: Data Repository Index
@@
Time Complexity.  Additionally, we provide a theoretical analysis of the time complexity for constructing the unified index, which is $O(mn\log\left(n\right))$, where $m$ and $n$ represent the number of data points in the dataset and the number of datasets in the data repository, respectively. The detailed analysis is provided in Appendix 9.2.
@@
6. Optimization of Search Layer
Based on the unified index constructed in the previous section, we propose effective pruning mechanisms to solve the data search problem defined in Section 3.5.
6.1. Accelerating Dataset Search
In this subsection, we describe the dataset search algorithm applied to range-based dataset search and top-$k$ exemplar dataset search defined in Section 3.3. Algorithm 2 shows the details of the implementation of the algorithm.
6.1.1. Accelerating Range-based Dataset Search
@@ 21                   $dis\leftarrow$ compute $dis$ between $Q$ and $N_{c}$;
6.1.2. Accelerating Top-$k$ Exemplar Dataset Search
@@
Figure 7. The distance bounds between two nodes using bounding boxes (left) and bounding balls (right).
Fast Bound Estimation.  We propose an efficient algorithm based on bounding balls of nodes to estimate the lower and upper bounds of the Hausdorff distance between nodes, which is not affected by the dimensions and only needs a single distance computation between center points. In contrast to the R-tree based incremental approach (Nutanong2011, ), which requires enumerating all pairs of corners of nodes’ bounding boxes (as illustrated by the four black dotted lines in Figure 7(a)), our method avoids quartic Euclidean distance computations (Nutanong2011, ) and only requires a single computation.
@@
(4) $\small\left[\max(||o_{1},o_{2}||-r_{2},0),\sqrt{||o_{1},o_{2}||^{2}+r_{2}^{2}}+r_{1}\right].$
@@ 34            $minUB\leftarrow\min(minUB,\textsf{UpperBound}(N_{Q},N_{C}))$;
Exact Hausdorff.  The above bounds can also be directly used in the fast pairwise Hausdorff distance computation. The main idea is to rank candidates based on two priority queues. Algorithm 3 shows how to prune distance computations based on bounds and achieve early termination without computing the full distance between all pairs of points from $Q$ and $D$. Similar to the incremental traversal (Please refer to Algorithm 4 in (Nutanong2011, )) [^2] Readers can refer to Algorithm 4 of (Nutanong2011, ) for more details. Here, we will skip them as the estimation of the bounds is our main focus., since the goal of Hausdorff is to calculate the maximum nearest neighbor distance, we first create a priority queue in descending order, ranking the distance from the point to their nearest neighbors. Then to compute the minimum distances to the nearest neighbors for points in $Q$, we further create an ascending priority queue to traverse the dataset from close to far. Subsequently, we traverse the index of $Q$ and $D$ from the root nodes, estimate the bounds between the polled pair of nodes, and conduct pruning. The whole algorithm terminates when the first pair of points are polled from the main priority queue. For convenience, we refer to this method as “ExactAppro” in Section 7.
Approximate Hausdorff.  Computing the exact Hausdorff distance can be time-consuming by going to the lowest point level of the dataset index. Instead, we can calculate an error-bounded Hausdorff distance in a more efficient way based on the unified index. In the exact distance computation, we terminate until a pair of points show up from the queue. To avoid such intensive distance computations, we set a rule that once a pair of “small” nodes have a radius smaller than a given threshold $\epsilon$ (see Line 3), we compute their Euclidean distance between their center points. Then we return the distance $||o_{1},o_{2}||$ as the approximate distance, and do not enqueue their child nodes or points anymore. For convenience, we refer to this method as “ApproHaus” in Section 7. Next, we prove that the ApproHaus method in Lemma 1 achieves a guaranteed precision of $2\epsilon$, where $\epsilon$ is the error threshold defined by the user.
@@
(8) $\epsilon=\frac{b^{\uparrow}[0]-b^{\downarrow}[0]}{2^{\theta}}.$
6.2. Accelerating Data Point Search
6.2.1. Range-based Point Search
@@
6.2.2. Nearest Neighbor Point Search
@@
Novelty of Acceleration Strategies.  According to the links between these technological developments, we can see that the proposed acceleration strategies based on our two-level unified index can solve multiple types of search problems at the same time compared to previous methods, which can only solve a single search problem (LiFX12, ; Guttman1984, ; Bentley75, ). In addition, we theoretically derive Hausdorff’s upper and lower bounds using the triangle inequality, and we explore the approximate Hausdorff value in depth with a guaranteed theoretical error.
7. Experiments
We conduct experiments on six real-world dataset repositories to demonstrate the effectiveness and efficiency of Spadas. In Section 7.1, we introduce the experimental setups. Then, we present the performance of our proposed unified index in Section 7.2. Next, we present the experimental results that vary with several key parameters to demonstrate the effectiveness and efficiency of the dataset search and data point search in Sections 7.3 and 7.4, followed by our case study in Section 7.5.
Table 1. Summary of six spatial data repositories.
@@ 7. Experiments
Table 2. Parameter settings.
@@
7.1. Settings
Datasets.  We conduct experiments on six representative spatial dataset repositories, each of which is downloaded from open-source spatial data portals. Table 1 presents the scale and dimension of each data repository. Figure 8 shows the distributions of the number of points in each data repository.
  • MultiOpen dataset, a real-world spatial data repository comprising datasets collected from three different open-source platforms: Points of Interests (POIs) of Baidu Maps Open Platform (baidu, ), BTAA Geoportal (btaa, ) and NYU Spatial Data Repository (nyu, ).
  • T-drive dataset (Yuan2010, ) contains a collection of real-world trajectory datasets generated by over 33,000 taxis in Beijing;
  • Argoverse self-driving datasets (ChangLS2019, ) which record five-second driving scenarios with thousands of points;
  • ShapeNet (Chang2015a, ), large-scale datasets of shapes represented by 3D computer-aided design (CAD) models of objects;
  • Chicago taxi trips (chicago, ) dataset, which has 194,992,061 trips in total. As all the trips are quantized every 15 minutes, we split the whole dataset every 15 minutes to create a set of sub-datasets, and there are 283,375 datasets in total;
  • Porto dataset (porto, ) transformed from 400+ taxis’ trajectories in the city of Porto for one year.
Figure 8. Dataset scale distributions in six data repositories.
Implementation.  We implement Spadas using Java 1.8 and run all our experiments on a 2x Intel Xeon Platinum 8268 24C 205W 2.9GHz Processor with 32G memory. Our code and guidelines for reproducibility can be accessed on the GitHub (spadas-code, ). We also have an online demonstration system (spade-system, ) on the Argoverse data repository, which supports the dataset and data point search proposed in this paper, and our proposed top-$k$ EMD search (YangWZ2022, ) is also available.
@@ 7.1. Settings
Parameter Settings.  The settings of parameters are summarized in Table 2 where the default settings are underlined. Below is the description of each parameter:
@@
  • $R$: It determines the range of query rectangular, where $\epsilon$ is the value obtained when the resolution $\theta$ is the default value $5$ according to the Equation 8 in Section 6.1.2.
@@
  • ScanGBO (Peng2016, ): the baseline comparison method needs to scan the candidate datasets to compute the grid-based overlap for each query dataset (see Algorithm 2 in (Peng2016, )).
  • ScanHaus (Adelfio2011, ): It designs the lower and upper bounds of Hausdorff by using two point sets’ MBR, and combines branch-and-bound to accelerate filter out dissimilar datasets.
  • IncHaus (Nutanong2011, ): It speeds up the Hausdorff distance computation by incrementally exploring the index of two point sets.
  • INNE (bandaragoda2018isolation, ; zhao2019pyod, ): the isolation-based anomaly detection using nearest-neighbor ensembles for outlier removal.
@@
  • kNN (Taha2015, ): a nearest neighbor search method using the early break strategy.
7.2. An Overview of spadas
We first show the running time of seven main steps in Spadas system, including the unified index construction (Index), dataset search (RangeS, IA, GBO, ExactHaus), and data point search (RangeP, NNP). As shown in Figure 9, we can observe that unified index construction is the most time-consuming step in most cases. However, the time taken to construct the index is still less than one minute. Moreover, the top-$k$ Hausdorff search is the slowest in dataset search compared with the IA and GBO due to its high complexity, but its running time is still tolerable.
Efficiency of the Unified Index.  We compare our unified index with the existing work that constructs an independent index using R-tree (Nutanong2011, ) in IncHaus. Figure 10 presents the construction cost and the occupied space as the data repository scale $m$ increases. It is evident that our proposed unified index outperforms the state-of-the-art method in terms of both index construction time and space footprint. Specifically, with regard to index construction time, the unified index can achieve at most 36.8% speedup over the index of IncHaus. Additionally, in terms of index space, the unified index is reduced by up to 90% over the IncHaus.
Figure 9. Efficiency in all stages on various data repositories.
Figure 10. Index’s construction time and occupied space with increasing data repository scale $m=\{10^{-3},10^{-2},10^{-1}$, 1}×$\times|\mathcal{D}|$.
Figure 11. Top-$k$ overlap-based queries as $k$ increases.
Figure 12. Top-$k$ overlap-based queries as $f$ increases.
Figure 13. Grid resolution $\theta$’s effects on top-$k$ GBO query.
Figure 14. Top-$k$ exact Hausdorff search with increasing $k$.
Figure 15. Top-k approximate Hausdorff search with increasing threshold $\epsilon$ (i.e., decreasing resolution $\theta$ in x-axis).
7.3. Efficiency of Dataset Search